# INT-PUB-25-009, LA-UR-25-23551, PSI-PR-25-07, YITP-SB-2025-08, ZU-TH 23/25 Baryon Number Violation: From Nuclear Matrix Elements to BSM Physics

### v1

@@ 5.5 The HIBEAM and NNBAR projects at the European Spallation Source
A conceptual design for a new 200 m beamline and experiment NNBAR has been carried out [8]. This would make use of a new large beam port and ideally a second moderator under the present ESS target wheel. Detailed MCNP and GEANT4 simulations of beamline and detector demonstrate that a sensitivity to observation of $n\rightarrow\bar{n}$ 1000$\sim$ 1000$\times$ 1000 × greater than ILL can be achieved. This clearly shows the potential for ESS to explore new regions of parameter space, but priorities and funding remain constrained in the short term. The chosen way forward is therefore to pursue a new, shorter particle physics beamline in the existing East experimental hall. The High-Intensity Baryon Extraction and Measurement (HIBEAM) collaboration is now designing this beamline and related infrastructure. The HIBEAM detector will make use of the existing WASA [67] crystal calorimeter to keep costs down. Detailed modeling shows the potential to go a factor of 10 beyond ILL in discovery reach (a factor of $\sim 3$ in lifetime sensitivity) with 3–5 years of running at 5 MW, as shown in Fig. 2. HIBEAM can also explore other physics including searches for sterile neutrons and axion dark matter.
@@ 5.6 A bottom-up approach to nucleon decay: RGEs, correlations, and connection to UV
[…]
@@ 5.7 Two-loop anomalous dimensions of $\Delta B$ operators
In this talk, we explain an on-going effort towards a full determination two-loop $\Delta B$ RGE at dimension six, paving the way for comprehensive future EFT analyses. Even though the considered sector does not lead to problems related to $\gamma_{5}$, the two-loop RGE is still scheme dependent. We discuss scheme dependencies that arise in $\Delta B$ sectors and we provide two well-defined schemes to be used: one in naive dimensional regularization and a second one in the ’t Hooft–Veltman scheme for $\gamma_{5}$ [85]. Our schemes feature advantages such as: restoration of spurion chiral symmetry and proper definition and handling of evanescent operators [86]. We also discuss a practical method for the calculation of the desired two-loop RGE. The method uses the well-known $R$ operation, which constructs all counterterms automatically, but must be adjusted to respected our schemes. The method has already proven useful in higher-loop EFT renormalization calculations [33]. We find preliminary results that pass all consistency checks such as: gauge invariance, pole relations, and restoration of spurious chiral symmetry. The results are part of a broader community effort towards two-loop anomalous dimensions and, more generally, next-to-leading logarithmic EFT analyses [87, 88, 89, 90].
@@ 5.9 Indirect constraints on the third-generation baryon number violation
[…]
@@ 5.11 Towards baryon number violation in nuclei from lattice QCD
$n\bar{n}$ oscillations, which violate $B$ and $B-L$, are predicted as low-energy signatures of BSM post-sphaleron baryogenesis scenarios [119, 120]. Experimental bounds on the $n\bar{n}$ transition time, $\tau_{n\bar{n}}$, have been obtained from searches for free neutron transitions [20] as well as intranuclear $n\bar{n}$ [121, 19]. New discoveries or limits will be set for free neutrons at the ESS [122] and for intranuclear $n\bar{n}$ by Hyper-Kamiokande and DUNE [123].
Relating $\tau_{n\bar{n}}$ to BSM parameters is accomplished in SMEFT by introducing a complete basis of six-quark operators, $Q_{I}$ [92]. Predictions for particular BSM theories are encoded in the Wilson coefficients $C_{I}$ of the Hamiltonian $\mathcal{H}_{n\bar{n}}=\sum_{I}C_{I}Q_{I}$. Its matrix element gives the $n\bar{n}$ transition rate $\tau_{n\bar{n}}^{-1}=\left|\left<\bar{n}|\mathcal{H}_{n\bar{n}}|n\right>\right|$. LQCD has been used to compute the $\bar{n}|Q_{I}|n\right>$ in a convenient basis [97] where the $Q_{I}$ renormalize multiplicatively and there are only four non-zero matrix elements [124, 125]. For example, in the left-right symmetric model of Ref. [126], the ILL bound and these LQCD results provide a bound of $350(22)$ TeV on the scale of $B$ violation.
@@
The simplest multi-nucleon matrix element of $Q_{I}$ for LQCD is $0|Q_{I}|nn\right>$, with $nn$ the two-neutron ground state. The corresponding quark-line diagram is simply the crossed version of the diagram describing $n\bar{n}$. Further, this matrix element can be identified as an “overlap factor” in a two-point correlation function involving $Q_{I}$ and an operator for the $nn$ ground state. Conveniently, such two-point functions were recently computed in variational studies of the $nn$ spectrum by the NPLQCD Collaboration in Ref. [127]. The $Q_{I}$ operator does not appear to couple strongly to the $nn$ ground state and has significantly larger overlap with the states coupling to plane-wave product operators $n(\mathbf{k})n(-\mathbf{k})$ with non-zero $\mathbf{k}$. This suggests that intranuclear $n\bar{n}$ rates could have significant dependence of the neutron momentum distributions of nuclei.
Current LQCD results only constrain the bare $0|Q_{I}|nn\right>$, and quantitative results must wait for determinations of the corresponding renormalization factors. Further challenges associated with energy-eigenstate identification for multi-nucleon systems can be mitigated using a new analysis framework based on the Lanczos algorithm that can provide rigorous two-sided bounds on the systematic errors associated with excited-state contamination [128, 129, 130]. Complete LQCD determinations of $0|Q_{I}|nn\right>$, for two-neutron ground and excited states, could then be matched to chiral EFT to constrain the relation between $\Gamma_{d}^{-1}$ and $\tau_{n\bar{n}}$. After matching LQCD and chiral EFT in the two-nucleon sector, the same relations would inform chiral-EFT predictions relating intranuclear $n\bar{n}$ rates to $\tau_{n\bar{n}}$ in larger nuclei.
@@ 5.12 Lattice QCD computations of proton decay matrix elements
Necessary steps for computing the matrix element in a certain renormalization scheme and scale are 1) computation of the matrix elements from three-point functions and 2) renormalization and matching of the BNV three-quark operators. There is a well-established way for 1), first done by JLQCD [112]. The difference from that to the current state-of-the-art computation is the use of dynamical fermion ensembles generated for the physical average up- and down-quark mass and physical strange-quark mass. 2) is done using the Rome–Southampton, RI/MOM (regularization independent, momentum subtraction) nonperturbative renormalization scheme constructed for proton-decay operators [132]. The RI/MOM scheme acts as an intermediate renormalization to bridge the lattice results to more convenient renormalization schemes like $\overline{\text{MS}}$, and is designed to absorb the lattice discretization details. The rest, matching to, e.g., $\overline{\text{MS}}$, can be done in continuum perturbation theory.
There are several different RI/MOM schemes so far used. For the quark wavefunction renormalization we typically utilize the vertex functions of bilinear operators with vanishing anomalous dimension, such as vector and axial vector operators. To avoid contamination of IR physics effect, RI/SMOM schemes [133, 134, 135] making use of non-exceptional momenta are applied for the bilinear vertex functions.
@@ 5.15 Baryon violation in cosmic rays: UHECR and cosmic antinuclei
Dark matter in the Universe may exist in the form of mirror (M) matter represented by the parallel hidden sector of particles which is an exact replica of ordinary (O) particle sector. Then all O particles: electron $e$, proton $p$, neutron $n$, neutrinos $\nu$, etc. must have their M twins $e^{\prime},p^{\prime},n^{\prime},\nu^{\prime}$ etc., which are sterile to the SM interactions, but have their own SM gauge interactions (for a review, see Ref. [151]). Cosmological bounds require M sector to be colder than O sector, $x=T^{\prime}/T\ll 1$, which can be realized by asymmetric reheating [152] and by demanding the out-of-equilibrium condition on cross-interactions between the O and M particle species at post-inflation epoch [153]. Namely, M baryons pass the CMB and LSS tests if $x<0.2$ or so [154]. In M sector, being colder and helium-dominated, stars are formed earlier and evolve faster [155], and their collapse can produce heavy black holes which may dominate galactic halos along with M stars in solar mass range.
@@
Here I discuss recent astrophysical hints which would be hard to explain without invoking $n$$n^{\prime}$ mixing. These are the following: (A) couple of years ago S. Ting reported that AMS-2 experiment has observed 8 candidate events for anti-helium nuclei in cosmic rays. Discovery of a single anti-nucleus with $A\geq 4$, if confirmed definitely, would challenge all known physics and cosmology, and its impact can be comparable to that of the positron discovery in 1932. (B) paradoxical situation in the UHECR: cosmic ray fluxes detected by two giant experiments, Pierre Auger Observatory (PAO) mostly exposed to South Hemisphere and Telescope Array (TA) exposed to North Hemisphere, are perfectly compatible at energies below the GZK cutoff $E_{\text{GZK}}\simeq 50$ EeV but become discrepant at higher energies, most notably at $E>100$ EeV. In addition, since the propagation length of such energetic UHECR is typically $\ell\sim$ 10 Mpc, they should come from potential sources closer than 30 Mpc or so. However, their arrival directions have no correlations with the nearby sources, and it seems that they rather come from the most empty regions of the sky. In particular, extreme energy event ($E=244$ EeV) recently detected by TA collaboration [170] points towards the center of the Local Void.
@@
(B) As far as M sector is colder, $x=T^{\prime}/T<0.2$ or so, the UHECR from M sources practically do not suffer the GZK cutoff and their flux can be 2–3 orders of magnitude larger than that of ordinary ones, since their propagation length scales as $\ell^{\prime}\simeq\ell/x^{3}$ and it can be as large as 1 Gpc or so [165]. Nevertheless, mirror UHECR with energies $E>x^{-1}E_{\text{GZK}}$, in their collisions with mirror relic photons, can produce the M neutrons $n^{\prime}$ which, provided that $\tau_{\epsilon}\ll\tau_{n}$, can promptly oscillate into ordinary neutrons $n$ and then produce protons via $\beta$ decay $n\to pe\bar{\nu}$. The energetic protons produced in this way within the ordinary GZK radius (i.e., in the nearby sky at distances $d<30$ Mpc or so), can safely reach us [166]. Interestingly, the condition for $n^{\prime}$$n$ oscillation not be effective reads as $(\tau_{\epsilon}/1\,\text{s})(B/1~{}\text{fG})<1$, and it can be satisfied only in cosmic voids where magnetic fields can be as small as $B<1$ fG. This would naturally explain why the most of extreme energy events come from the voids. In addition, since the Northern Sky is more “voidy,” and namely the Local Void is mostly exposed to the TA experiment, this can also explain the difference between the PAO and TA fluxes at $E>100$ EeV.
Summarizing, the situation looks very optimistic for the ongoing and planned searches at different neutron facilities at the PSI, ILL, ORNL, and ESS [172, 173, 174]. The NS bounds on $n$$n^{\prime}$ conversion imply $\tau_{\epsilon}>1$ s or so [167], and in this case production of antimatter via $\bar{n}^{\prime}\to\bar{n}$ in the cores of mirror NS can explain the paradox (A) on the cosmic anti-helium flux. On the other hand, solution of the paradox (B) quests for $\tau_{\epsilon}<100$ s or so. This gives a hope that under the systematic search these experiments may indeed discover $n$$n^{\prime}$ oscillation in the range $\tau_{\epsilon}=1\div 100$ s.
@@ 5.16 Some recent results on $n\bar{n}$ transitions
The violation of baryon number is one of the necessary conditions for baryogenesis. In addition to proton decay, an interesting possibility for BNV is $n\bar{n}$ transitions [126]. Although the SM conserves $B$ perturbatively, BNV is present in many BSM theories. In this talk we first review the basic formalism describing $n\bar{n}$ transitions in quasi-free neutron propagation experiments and in matter, and then discuss a model in which proton decay can be suppressed well below current experimental limits while $n\bar{n}$ transitions and associated $\Delta B=-2$ dineutron decays can occur near to current limits [175, 176, 177]. Thus, in this model, $n\bar{n}$ transitions and associated $\Delta B=-2$ dineutron decays are the main manifestation of BNV, rather than proton decay. This model features a certain number of extra compactified spatial dimensions, with left- and right-handed chiral components of SM fermion fields having wave functions that are strongly localized at various different points in the extra dimensions. It has the appeal that it can naturally explain the generational hierarchy in quark and charged lepton masses [178], and can also explain the masses and mixings of neutrinos [179]. The effective 4D Lagrangian and coefficients of six-quark operators responsible for the $n\bar{n}$ transitions are calculated by integrating over the corresponding operators in the extra dimensions. This model provides further motivation for the future planned $n\bar{n}$ search experiment at the ESS [173], see also Refs. [180, 181]. Further, we discuss an analysis of the effects of $n\bar{n}$ transitions in neutron stars [182]. Each $n\bar{n}$ transition leads to $\Delta B=-2$ annihilation, releasing $\sim 2m_{n}$ energy. We calculate the effect on the luminosity of the neutron star and find that it is negligibly small. We also calculate the effects on the rotation of neutron stars and on the periods of binary pulsars, again finding these to be negligibly small.

### v2

@@ 5.5 The HIBEAM and NNBAR projects at the European Spallation Source
A conceptual design for a new 200 m beamline and experiment NNBAR has been carried out [8]. This would make use of a new large beam port and ideally a second moderator under the present ESS target wheel. Detailed MCNP and GEANT4 simulations of beamline and detector demonstrate that a sensitivity to observation of $n\rightarrow\bar{n}$ $\sim 1000\times$ greater than ILL can be achieved. This clearly shows the potential for ESS to explore new regions of parameter space, but priorities and funding remain constrained in the short term. The chosen way forward is therefore to pursue a new, shorter particle physics beamline in the existing East experimental hall. The High-Intensity Baryon Extraction and Measurement (HIBEAM) collaboration is now designing this beamline and related infrastructure. The HIBEAM detector will make use of the existing WASA [75] crystal calorimeter to keep costs down. Detailed modeling shows the potential to go a factor of 10 beyond ILL in discovery reach (a factor of $\sim 3$ in lifetime sensitivity) with 3–5 years of running at 5 MW, as shown in Fig. 2. HIBEAM can also explore other physics including searches for sterile neutrons and axion dark matter.
@@ 5.6 A bottom-up approach to nucleon decay: RGEs, correlations, and connection to UV
Motivated by the upcoming series of experiments searching for proton decay [6, 5, 66], this work focuses on BNV from a bottom-up perspective. Specifically, we characterize nucleon decay in a model-independent framework based on SMEFT [76, 77, 78], highlighting the importance of searching for both $|\Delta(B-L)|=0,\;2$ nucleon decay modes. These modes are often indistinguishable from an experimental standpoint, yet their observation would point to different SMEFT operators, which originate from distinct UV models.
@@ 5.7 Two-loop anomalous dimensions of $\Delta B$ operators
Here, we explain an on-going effort towards a full determination two-loop $\Delta B$ RGE at dimension six, paving the way for comprehensive future EFT analyses. Even though the considered sector does not lead to problems related to $\gamma_{5}$, the two-loop RGE is still scheme dependent. We discuss scheme dependencies that arise in $\Delta B$ sectors and we provide two well-defined schemes to be used: one in naive dimensional regularization and a second one in the ’t Hooft–Veltman scheme for $\gamma_{5}$ [93]. Our schemes feature advantages such as: restoration of spurion chiral symmetry and proper definition and handling of evanescent operators [94]. We also discuss a practical method for the calculation of the desired two-loop RGE. The method uses the well-known $R$ operation, which constructs all counterterms automatically, but must be adjusted to respected our schemes. The method has already proven useful in higher-loop EFT renormalization calculations [33]. We find results [95] that pass all consistency checks such as: gauge invariance, pole relations, and restoration of spurious chiral symmetry. The results are part of a broader community effort towards two-loop anomalous dimensions and, more generally, next-to-leading logarithmic EFT analyses [96, 97, 98, 99].
@@ 5.9 Indirect constraints on the third-generation baryon number violation
$\mathcal{H}_{p\to\ell^{+}\nu_{\ell}\bar{\nu}}$ $=-2\sqrt{2}\,\frac{G_{F}C_{\nu}V_{ub}^{*}}{m_{b}\Lambda_{\rm BNV}^{2}}\,\mathcal{O}_{\nu,\text{sl}}+\text{h.c.}\,$,
$\mathcal{H}_{p\to\pi^{+}\bar{\nu}}$ $=-2\sqrt{2}\,\frac{G_{F}C_{\nu}V_{ub}^{*}V_{ud}}{m_{b}\Lambda_{\rm BNV}^{2}}\Bigl{(}C_{1}\mathcal{O}_{\nu,1}+C_{2}\mathcal{O}_{\nu,2}\Bigr{)}+\text{h.c.}\,$,
$\mathcal{H}_{p\to\pi^{0}\ell^{+}}$ $=-2\sqrt{2}\,\frac{G_{F}V_{ub}^{*}V_{ud}}{m_{b}\Lambda_{\rm BNV}^{2}}\sum_{X=L,R}C_{X}\Bigl{(}C_{1}\mathcal{O}_{X,1}+C_{2}\mathcal{O}_{X,2}\Bigr{)}+\text{h.c.}\,$, (1)
@@ 5.11 Towards baryon number violation in nuclei from lattice QCD
$n\bar{n}$ oscillations, which violate $B$ and BLB-L, are predicted as low-energy signatures of BSM post-sphaleron baryogenesis scenarios [128, 129]. Experimental bounds on the $n\bar{n}$ transition time, $\tau_{n\bar{n}}$, have been obtained from searches for free neutron transitions [20] as well as intranuclear $n\bar{n}$ [61, 19]. New discoveries or limits will be set for free neutrons at the ESS [130] and for intranuclear $n\bar{n}$ by Hyper-Kamiokande and DUNE [131].
[…]
@@
The simplest multi-nucleon matrix element of $Q_{I}$ for LQCD is $\left<0|Q_{I}|nn\right$>, with $nn$ being the two-neutron ground state. The corresponding quark-line diagram is simply the crossed version of the diagram describing $n\bar{n}$. Further, this matrix element can be identified as an “overlap factor” in a two-point correlation function involving $Q_{I}$ and an operator for the $nn$ ground state. Conveniently, such two-point functions were recently computed in variational studies of the $nn$ spectrum by the NPLQCD Collaboration in Ref. [135]. The $Q_{I}$ operator does not appear to couple strongly to the $nn$ ground state and has significantly larger overlap with the states coupling to plane-wave product operators $n(\mathbf{k})n(-\mathbf{k})$ with non-zero $\mathbf{k}$. This suggests that intranuclear $n\bar{n}$ rates could have significant dependence of the neutron momentum distributions of nuclei.
Current LQCD results only constrain the bare $\left<0|Q_{I}|nn\right$>, and quantitative results must wait for determinations of the corresponding renormalization factors. Further challenges associated with energy-eigenstate identification for multi-nucleon systems can be mitigated using a new analysis framework based on the Lanczos algorithm that can provide rigorous two-sided bounds on the systematic errors associated with excited-state contamination [136, 137, 138]. Complete LQCD determinations of $\left<0|Q_{I}|nn\right$>, for two-neutron ground and excited states, could then be matched to Chiral EFT to constrain the relation between $\Gamma_{d}^{-1}$ and $\tau_{n\bar{n}}$. After matching LQCD and Chiral EFT in the two-nucleon sector, the same relations would inform Chiral-EFT predictions relating intranuclear $n\bar{n}$ rates to $\tau_{n\bar{n}}$ in larger nuclei.
@@ 5.12 Lattice QCD computations of proton decay matrix elements
Necessary steps for computing the matrix element in a certain renormalization scheme and scale are 1) computation of the matrix elements from three-point functions and 2) renormalization and matching of the BNV three-quark operators. There is a well-established way for 1), first done by JLQCD [121]. The difference from that to the current state-of-the-art computation is the use of dynamical fermion ensembles generated for the physical average up- and down-quark mass and physical strange-quark mass. 2) is done using the Rome–Southampton, RI/MOM [140] (regularization independent, momentum subtraction) nonperturbative renormalization scheme constructed for proton-decay operators [141]. The RI/MOM scheme acts as an intermediate renormalization to bridge the lattice results to more convenient renormalization schemes like $\overline{\text{MS}}$, and is designed to absorb the lattice discretization details. The rest, matching to, e.g., $\overline{\text{MS}}$, can be done in continuum perturbation theory.
There are several different RI/MOM schemes used so far. For the quark wavefunction renormalization we typically utilize the vertex functions of bilinear operators with vanishing anomalous dimension, such as vector and axial vector operators. To avoid contamination of IR physics effect, RI/SMOM schemes [142, 143, 144] making use of non-exceptional momenta are applied for the bilinear vertex functions.
@@ 5.15 Baryon violation in cosmic rays: UHECR and cosmic antinuclei
Dark matter in the Universe may exist in the form of mirror (M) matter represented by the parallel hidden sector of particles which is an exact replica of ordinary (O) particle sector. Then all O particles: electron $e$, proton $p$, neutron $n$, neutrinos $\nu$, etc. must have their M twins $e^{\prime},p^{\prime},n^{\prime},\nu^{\prime}$ etc., which are sterile to the SM interactions, but have their own SM gauge interactions (for a review, see Ref. [160]). Cosmological bounds require M sector to be colder than O sector, $x=T^{\prime}/T\ll 1$, which can be realized by asymmetric reheating [161] and by demanding the out-of-equilibrium condition on cross-interactions between the O and M particle species at post-inflation epoch [162]. Namely, M baryons pass the cosmic-microwave-background and large-scale-structure tests if $x<0.2$ or so [163]. In M sector, being colder and helium-dominated, stars are formed earlier and evolve faster [164], and their collapse can produce heavy black holes which may dominate galactic halos along with M stars in solar mass range.
@@
Here I discuss recent astrophysical hints which would be hard to explain without invoking $nn^{\prime}$ mixing. These are the following: (A) two years ago it was reported that the AMS-2 experiment had observed 10 candidate events for anti-helium nuclei in cosmic rays [179]. Discovery of a single anti-nucleus with $A\geq 4$, if confirmed definitely, would challenge all known physics and cosmology, and its impact can be comparable to that of the positron discovery in 1932. (B) paradoxical situation in the UHECR: cosmic ray fluxes detected by two giant experiments, Pierre Auger Observatory (PAO) mostly exposed to South Hemisphere and Telescope Array (TA) exposed to North Hemisphere, are perfectly compatible at energies below the GZK cutoff $E_{\text{GZK}}\simeq 50$ EeV but become discrepant at higher energies, most notably at $E>100$ EeV. In addition, since the propagation length of such energetic UHECR is typically $\ell\sim$ 10 Mpc, they should come from potential sources closer than 30 Mpc or so. However, their arrival directions have no correlations with the nearby sources, and it seems that they rather come from the most empty regions of the sky. In particular, the extreme energy event ($E=244$ EeV) recently detected by TA collaboration [180] points towards the center of the Local Void.
@@
(B) As far as M sector is colder, $x=T^{\prime}/T<0.2$ or so, the UHECR from M sources practically do not suffer the GZK cutoff and their flux can be 2–3 orders of magnitude larger than that of ordinary ones, since their propagation length scales as $\ell^{\prime}\simeq\ell/x^{3}$ and it can be as large as 1 Gpc or so [174]. Nevertheless, mirror UHECR with energies $E>x^{-1}E_{\text{GZK}}$, in their collisions with mirror relic photons, can produce the M neutrons $n^{\prime}$ which, provided that $\tau_{\epsilon}\ll\tau_{n}$, can promptly oscillate into ordinary neutrons $n$ and then produce protons via $\beta$ decay $n\to pe\bar{\nu}$. The energetic protons produced in this way within the ordinary GZK radius (i.e., in the nearby sky at distances $d<30$ Mpc or so), can safely reach us [175]. Interestingly, the condition for $^{\prime}nn$ oscillation to not be effective reads as $(\tau_{\epsilon}/1\,\text{s})(B/1~\text{fG})<1$, and it can be satisfied only in cosmic voids where magnetic fields can be as small as $B<1$ fG. This would naturally explain why the most of extreme energy events come from the voids. In addition, since the Northern Sky is more “voidy,” and namely the Local Void is mostly exposed to the TA experiment, this can also explain the difference between the PAO and TA fluxes at $E>100$ EeV.
Summarizing, the situation looks very optimistic for the ongoing and planned searches at different neutron facilities at the PSI, ILL, ORNL, and ESS [182, 183, 184]. The NS bounds on $nn^{\prime}$ conversion imply $\tau_{\epsilon}>1$ s or so [176], and in this case production of antimatter via $\bar{n}^{\prime}\to\bar{n}$ in the cores of mirror NS can explain the paradox (A) on the cosmic anti-helium flux. On the other hand, solution of the paradox (B) quests for $\tau_{\epsilon}<100$ s or so. This gives a hope that under the systematic search these experiments may indeed discover $nn^{\prime}$ oscillation in the range $\tau_{\epsilon}=(1\text{--}100)$s.
@@ 5.16 Some recent results on $n\bar{n}$ transitions
The violation of baryon number is one of the necessary conditions for baryogenesis. In addition to proton decay, an interesting possibility for BNV is $n\bar{n}$ transitions [134]. Although the SM conserves $B$ perturbatively, BNV is present in many BSM theories. Here, we first review the basic formalism describing $n\bar{n}$ transitions in quasi-free neutron propagation experiments and in matter, and then discuss a model in which proton decay can be suppressed well below current experimental limits while $n\bar{n}$ transitions and associated $\Delta B=-2$ dineutron decays can occur near to current limits [185, 186, 187]. Thus, in this model, $n\bar{n}$ transitions and associated $\Delta B=-2$ dineutron decays are the main manifestation of BNV, rather than proton decay. This model features a certain number of extra compactified spatial dimensions, with left- and right-handed chiral components of SM fermion fields having wave functions that are strongly localized at various different points in the extra dimensions. It has the appeal that it can naturally explain the generational hierarchy in quark and charged lepton masses [188], and can also explain the masses and mixings of neutrinos [189]. The effective 4D Lagrangian and coefficients of six-quark operators responsible for the $n\bar{n}$ transitions are calculated by integrating over the corresponding operators in the extra dimensions. This model provides further motivation for the future planned $n\bar{n}$ search experiment at the ESS [183], see also Refs. [190, 191]. Further, we discuss an analysis of the effects of $n\bar{n}$ transitions in neutron stars [58]. Each $n\bar{n}$ transition leads to $\Delta B=-2$ annihilation, releasing $\sim 2m_{n}$ energy. We calculate the effect on the luminosity of the neutron star and find that it is negligibly small. We also calculate the effects on the rotation of neutron stars and on the periods of binary pulsars, again finding these to be negligibly small.